Grade 12Physical Chemistry

Two liquids A and B form an ideal solution at temperature T. When the total vapour pressure above the solution is 400 torr, the mole fraction of A in the vapour phase is 0.40 and in the liquid phase 0.75. What are the vapour pressure of pure Aand pure B at temperature T?

Kevin woods
9 Years agoGrade 12

1 Answer

Khimraj
9 Years ago
As total pressure in liquid phase P = (PoA)(Xa) + (PoB)(Xb) so 400 = 3/4PoA + 1/4PoA. And in vapour phase 4/10 = Pa/(Pa+Pb) = (3/4PoA)/((3/4PoA)+(1/4PoB)). Now solve these equation and find PoA and PoB.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer